OpenStudy (anonymous):

Solving quadratics by factoring, solve equaton -20x^2+14=-27x

OpenStudy (anonymous):

first add 27x to each side -20x^2 +27x + 14 = 0

OpenStudy (anonymous):

-7x^2+1=0

OpenStudy (anonymous):

oh no, you can't subtract -20x^2 and 27x one is x to the second power

OpenStudy (anonymous):

-20x^2-13x=0

OpenStudy (anonymous):

it's -20x^2 + 27x + 14 = 0 the 27 has a variable of x which we don't know what is so we can't add or subtract with it unless another number has the same variable

OpenStudy (anonymous):

this algbra is confusing

OpenStudy (anonymous):

ok, so if your adding or subtracting with a variable you need to have the same variable, but if your multiply or dividing it doesn't matter.

OpenStudy (anonymous):

Thanks for ur help

OpenStudy (anonymous):

ok well we have -20x^2 +27x + 14 = 0 and ima pass it to @JakeV8 mainly because i forget this part, and i need sleep lol

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

ok Jake from here i need to find the smaller value & the larger value

OpenStudy (anonymous):

Sorry, I got kicked off the site and lost everything I had typed. This doesn't factor into something easy with integer values... -20x^2 +27x + 14 = 0 you can divide it all through by -20 to get: x^2 - (27/20)x + (14/20) = 0 But in my opinion, that's not much better. The solutions are x = -8/20 and x = 35/20 which both reduce to x = -2/5 and x = 7/4 that means you could have factored it as: (x + 2/5)(x - 7/4) = 0 If you can see how to easily get from the original quadratic equation to this factored form, then great. Personally, it doesn't seem easy or obvious to me, and if the problem hadn't stated to solve by factoring, I would have used the quadratic formula to get the solutions for x.